# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ 4.3.2 Regret of mirror descent policy optimisation with robust Lagrangian
The proof follows from the definition
$\Phi(\pi_{k})-\Phi(\pi^{*})$ $=\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi_{k},p}(\rho)-\sup_{p^{\prime}\in\mathcal{P}}\mathbf{V}_{\pi^{*},p^{\prime}}(\rho)$
[…]
$\leq\mathbf{V}_{\pi_{k},p_{k}}(\rho)+\epsilon_{k}^{\prime}-\mathbf{V}_{\pi^{*},p_{k}}(\rho)\,.$
We now turn to the theorem, which matches Wang and Petrik, (2024) up to a constant related to the range of the Lagrangian. The additional factor $\frac{1}{1-\gamma}\left\lVert\frac{1}{\mu}\right\rVert_{\infty}$ is consistent with the analysis of Mei et al., (2020), accounting for the possible mismatch between the initial sampling distribution $\mu$ and the true initial distribution $\rho$ of the MDP.
Theorem 1 (Regret of mirror descent policy optimisation with robust Lagrangian).
Using constant step size $\eta=1/l_{\theta}$, Lagrangian gradient descent over logits yields at any macro-step $k\geq 0$
$\Phi(\pi_{k})-\Phi(\pi^{*})\leq\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{32SM^{2}F_{\lambda_{k}}}{U(1-\gamma)^{6}k}+\epsilon_{k}^{\prime}\,,$ (27)
where $\epsilon_{k}^{\prime}>0$ is the tolerance of LTMA at macro-step $k$.
Proof.
Since for the softmax parametrisation, mirror descent is equivalent to traditional gradient descent (see Table 5), the robust Lagrangian policy updates will produce a sequence of parameters $\left\{\theta_{k}\right\}_{k=1}^{n}$ obtained from the process
$\theta_{k+1}=\theta_{k}-\eta\nabla\mathbf{V}_{\pi_{\theta_{k}},p_{k}}(\rho)\,.$ (28)
Due to the smoothness and continuity of the Lagrangian value in Lemma 5 and Lemma 4, it follows that for any $k^{\prime}\geq 1$, there exists a parameter sequence $\left\{\hat{\theta}_{k}\right\}_{k=1}^{k^{\prime}}$ obtained from non-robust policy gradient descent with nominal dynamics $\hat{p}=p_{k^{\prime}}$ and Lagrangian multiplier $\hat{\lambda}=\lambda_{k^{\prime}}$ such that $\hat{\theta}_{k^{\prime}}=\theta_{k^{\prime}}$, which is the following process with constant transition kernel:
$\hat{\theta}_{k+1}=\hat{\theta}_{k}-\eta\nabla\mathbf{V}_{\pi_{\hat{\theta}_{k}},\hat{p}}(\rho;\hat{\lambda})\,.$ (29)
Applying Theorem 6 and accounting for the worst-case dynamics by letting $U=\inf_{p\in\mathcal{P}}U_{p}$ and $M=\sup_{p\in\mathcal{P}}$, we obtain
$\mathbf{V}_{\pi_{k^{\prime}},p_{k^{\prime}}}(\rho)-\mathbf{V}_{\pi^{*},p_{k^{\prime}}}(\rho)\leq\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{32SM^{2}F_{\lambda_{k^{\prime}}}}{U(1-\gamma)^{6}k^{\prime}}\,.$ (30)
Combining with Eq. 26 from Lemma 7, we obtain
$\Phi(\pi_{k^{\prime}})-\Phi(\pi^{*})\leq\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{32SM^{2}F_{\lambda_{k^{\prime}}}}{U(1-\gamma)^{6}k^{\prime}}+\epsilon_{k^{\prime}}^{\prime}\,.$ (31)
Theorem 1 implies an $\mathcal{O}(T^{-1})$ convergence rate to the optimal robust Lagrangian.
With entropy regularisation
We now analyse the case where an additional negative entropy term is added to the Lagrangian. The result is useful to consider an additional term to the objective but also to consider the performance of mirror descent with KL-divergence based Bregman divergence. In the former case, the objective is formulated in terms of the entropy-regularised value function $\mathbf{V}_{\pi,p}^{\alpha}(\rho)-\tau\mathcal{H}(\rho,\pi,p)$ based on the discounted entropy
$\mathcal{H}(\rho,\pi,p)=-\mathbb{E}\left[\sum_{l=0}^{\infty}\gamma^{l}\log(\pi(a_{l}|s_{l}))|\pi,p\right]\,,$ (32)
such that deterministic policies are penalised more strongly. In the latter case, the objective is based on $\mathbf{V}_{\pi,p}(\rho)-\tau\mathcal{H}(\rho,\pi,p)$ but it is useful to note the interpretation of PMD-PD as the entropy-regularised objective with costs based on $\tilde{\mathbf{c}}_{k}(s,a)-\tau\log(\pi_{k}(a|s))$,
$\mathbf{V}_{\pi,p}^{\tau}(\rho)=\mathbb{E}\left[\sum_{l=0}^{\infty}\gamma^{l}\left(\underbrace{\mathbf{c}_{k}(s_{l},a_{l})+\tau\log\left(\frac{1}{\pi_{k}(a_{l}|s_{l})}\right)}_{\text{cost}}+\underbrace{\tau\log(\pi(a_{l}|s_{l}))}_{\text{entropy regularisation}}\right)|\pi,p\right]\,.$
The aforementioned properties of smoothness and continuity can be easily extended to the entropy-regularised Lagrangian.
Lemma 8 (Smooth and continuous entropy-regularised Lagrangian).
Let $\tau>0$, $\tau=\Theta(\frac{1}{\log(A)})$ and let $F_{\lambda}$ be chosen according to Eq. 11 or Eq. 12 (depending on the use of augmentation). Then under the softmax parametrisation, the entropy-regularised Lagrangian $\mathbf{V}(\theta)+\tau\mathcal{H}(\rho,\pi_{\theta})$ is $l_{\theta}$-smooth and $L_{\theta}$-Lipschitz continuous with
$l_{\theta}=\mathcal{O}(\frac{F_{\lambda}}{(1-\gamma)^{2}})$ (33)
and
$L_{\theta}=\mathcal{O}(\frac{F_{\lambda}}{(1-\gamma)^{3}})\,.$ (34)
Proof.
Let $k\geq 1$. From Lemma 5,
$|\mathbf{V}(\theta^{\prime})-\mathbf{V}(\theta)-\left\langle\nabla\mathbf{V}(\theta),\theta^{\prime}-\theta\right\rangle|\leq\frac{16F_{\lambda}}{(1-\gamma)^{3}}\quad\forall\theta,\theta^{\prime}\in\Theta\,.$
From Lemma 14 in Mei et al., (2020), it follows that
$|\mathcal{H}(\theta^{\prime})-\mathcal{H}(\theta)-\left\langle\nabla\mathcal{H}(\theta),\theta^{\prime}-\theta\right\rangle|\leq\frac{4+8\log(A)}{(1-\gamma)^{3}}\quad\forall\theta,\theta^{\prime}\in\Theta\,.$
Combining both, via triangle inequality we obtain
$|\mathbf{V}(\theta^{\prime})+\mathcal{H}(\theta^{\prime})-(\mathbf{V}(\theta^{\prime})+\mathcal{H}(\theta))-\left\langle\nabla(\mathbf{V}+\mathcal{H})(\theta),\theta^{\prime}-\theta\right\rangle|$ $\leq\frac{16F_{\lambda}+\tau\,(4+8\log(A))}{(1-\gamma)^{3}}\quad\forall\theta,\theta^{\prime}\in\Theta\,.$
It follows that $\mathbf{V}(\theta)+\mathcal{H}(\theta)$ is $l_{\theta}$-smooth with $l_{\theta}=\mathcal{O}(\frac{F_{\lambda}}{(1-\gamma)^{3}})$.
For Lipschitz continuity, a similar argument can be made. Note that since the entropy ranges in $[0,\log(A)]$, the Lipschitz bound on the discounted entropy is given by
$|\mathcal{H}(\theta^{\prime})-\mathcal{H}(\theta)|\leq\frac{\log(A)}{1-\gamma}\,.$
From derivations in Lemma 4.4 in Wang et al., (2023),
$\left\lVert\nabla\log(\pi(a|s))\right\rVert\leq\sqrt{2}\,.$
Denoting $\mathbf{V}^{\tau}$ as the regularised objective, we have
$\max_{\theta,\theta^{\prime}}\left\lVert\frac{\mathbf{V}^{\tau}(\theta)-\mathbf{V}^{\tau}(\theta^{\prime})}{\theta-\theta^{\prime}}\right\rVert$ $=\max_{\theta}\left\lVert\frac{\partial\mathbf{V}^{\tau}(\theta)}{\partial\theta(s,\cdot)}\right\rVert$
$\leq\frac{1}{(1-\gamma)}\left\lVert\nabla\log(\pi(a|s))\right\rVert\left\lVert\mathbf{V}^{\tau}(\theta)\right\rVert$
$\leq\frac{2\sqrt{2}(F_{\lambda}+\log(A))}{(1-\gamma)^{2}}$
$=\mathcal{O}\left(\frac{F_{\lambda}}{(1-\gamma)^{2}}\right)\,.$
The following theorem follows directly from Theorem 6 of Mei et al., (2020), allowing an exponential convergence rate under gradient descent with entropy regularisation.
Theorem 2 (Convergence rate for softmax policies for Lagrangian gradient descent with entropy regularisation).
Let $\Pi$ be the softmax parametrisation. For Lagrangian gradient descent over logits and the conditions of Lemma 8, with settings of $M=\inf_{p\in\mathcal{P},\pi\in\Pi}M_{p}(\pi)$, $U=\inf_{p\in\mathcal{P}}U_{p}$, and $\eta=\frac{1}{l_{\theta}}$ according to Theorem 5, it follows that for any $t\geq 0$, for a fixed transition kernel $p\in\mathcal{P}$, and fixed $\lambda_{k}=\lambda_{k-1}$,
$\mathbf{V}_{\pi_{t},p}^{\tau}(\rho)-\mathbf{V}_{\pi_{\tau}^{*},p}^{\tau}(\rho)=\mathcal{O}\left(\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{F_{\lambda}}{(1-\gamma)}e^{-C\,(t-1)}\right)\,,$ (35)
where $C=\frac{\eta}{S}\min_{s}\mu(s)U^{2}M^{-1}$ and $\pi_{\tau}^{*}\in\mathrm{arg}\max_{\pi\in\Pi}\mathbf{V}_{\pi,p}^{\tau}(\rho)$.
Finally, we apply Theorem 2 to the robust Lagrangian problem 4, or more specifically, to the robust entropy-regularised Lagrangian problem
@@
$\min_{\pi}\left\{\Phi^{\tau}(\pi):=\sup_{p\in\mathcal{P}}\max_{\lambda\geq 0}\mathbf{V}_{\pi,p}^{\tau}(\rho)\right\}\,.$ (36)
Theorem 3 (Regret of mirror descent policy optimisation with robust Lagrangian with entropy regularisation).
Using constant step size $\eta=1/l_{\theta}$, Lagrangian gradient descent over softmax logits with entropy regularisation yields at any macro-step $k\geq 0$
$\Phi^{\tau}(\pi_{k})-\Phi^{\tau}(\pi_{\tau}^{*})=\mathcal{O}\left(\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{F_{\lambda}}{(1-\gamma)^{2}}e^{-C\,(k-1)}+\epsilon_{k}^{\prime}\right)\,,$ (37)
where $F_{\lambda}$ is set according to Eq. 11 or Eq. 12 depending on the use of (un-)augmented Lagrangian, $\pi_{\tau}^{*}\in\mathrm{arg}\min_{\pi\in\Pi}\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi,p}^{\tau}(\rho)$, and $\epsilon_{k}^{\prime}>0$ is the tolerance of LTMA at macro-step $k$ with respect to $\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi_{,}p}^{\tau}(\rho)$.
Proof.
Analogous to the proof of Theorem 1, we bound the regret by
$\Phi^{\tau}(\pi_{k})-\Phi^{\tau}(\pi_{\tau}^{*})$ $\leq\mathbf{V}_{\pi_{\theta_{k}},p_{k}}^{\tau}(\rho)+\epsilon_{k}-\mathbf{V}_{\pi_{\tau}^{*},p_{k}}^{\tau}(\rho)\,,$
and formulate an equivalent gradient descent process
$\hat{\theta}_{k+1}=\hat{\theta}_{k}-\eta\nabla\mathbf{V}_{\pi_{\hat{\theta}_{k}},\hat{p}}^{\tau}(\rho;\hat{\lambda})\,.$ (38)
Applying Theorem 2, we obtain for any $k^{\prime}\geq 1$ that
$\Phi^{\tau}(\pi_{k^{\prime}})-\Phi^{\tau}(\pi_{\tau}^{*})=\mathcal{O}\left(\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{F_{\lambda}}{(1-\gamma)}e^{-C\,(k^{\prime}-1)}+\epsilon_{k^{\prime}}^{\prime}\right)\,.$ (39)
Theorem 2 implies an $\mathcal{O}(e^{-T})$ convergence rate to the optimal regularised robust Lagrangian, which is not necessarily optimal on the original robust Lagrangian problem.
4.4 Sample-based analysis
Having shown results for the oracle-based setting where values and policy gradients are given, we now turn to the sample-based analysis, where values and policy gradients need to be estimated from finite-sample data are accounted for in the convergence rate computations. The analysis modifies the techniques from Sample-based PMD-PD (Liu et al.,, 2021) by accounting for the mismatch between the nominal transition kernel and the optimal transition kernel in the min-max problem in Eq. 3. Their analysis makes use of a weighted Bregman divergence term formulated in Eq. 13. While the weighted Bregman divergence is not a proper Bregman divergence over policies, it is over the occupancy distribution, i.e. $B_{d_{\rho}^{\pi,p}}(\pi,\pi^{\prime})=B(d_{\rho}^{\pi,p},d_{\rho}^{\pi^{\prime},p})$, as shown in Lemma 10 of Liu et al., (2021). The complexity analysis below is slightly simplified by dropping purely problem-dependent constants $\gamma$, $\eta$, $\zeta$ from the big-O statements.
@@ 4.4 Sample-based analysis
As sample-based algorithms work with approximate value functions, Lemma 9 describes the conditions under which one can obtain an $\epsilon$-approximation of the cost functions and the regularised augmented Lagrangian. The same settings as sample-based PMD in Lemma 15 of Liu et al., (2021) transfer to the robust sample-based PMD algorithm. We rephrase the lemma to include updated transition dynamics.
Lemma 9 (Value function approximation, Lemma 15 of Liu et al., (2021) rephrased).
@@
[…]
@@
The full proof can be found in Appendix C.5. ∎
@@
Lemma 10 (Bound on regularised augmented Lagrangian under approximate entropy-regularised NPG, Lemma 18 in Liu et al., (2021)).
@@
$\tilde{\mathbf{V}}_{\pi_{k+1},p_{k}}(s)+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k+1},p_{k}}}(\pi_{k+1},\pi_{k})\leq\tilde{\mathbf{V}}_{\pi,p_{k}}(s)+\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k})-B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k+1})\right)+\Theta(\epsilon)$
@@
The proof makes use of Lemma 18. In particular, it notes that $K=\Theta(1/\epsilon)$ and the value difference to $\pi_{k}^{*}$ is bounded by $1/K$ after $t_{k}$ steps. After applying the pushback property (Lemma 15) to $\pi$, $\pi_{k}$, and $\pi_{k}^{*}$, and derives an upper bound on the Bregman divergence w.r.t. $\pi_{k}^{*}$ based on $B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k+1})+\left\lVert\log(\pi_{k}^{*})-\log(\pi_{k}^{t+1})\right\rVert_{\infty}$, where the second term is $\Theta(1/K)$. We refer to Lemma 18 of Liu et al., (2021) for the detailed proof. ∎
Below we prove the sample complexity of robust sample-based PMD-PD. For policy updates, we use estimates of the Q-values and show the overall sample complexity for obtaining an average regret in the value, constraint-cost, and Lagrangian bounded by $\mathcal{O}(\epsilon)$. Since the theory on TMA is not yet well developed, the analysis will assume the oracle setting for transition kernel updates, in which case we simply count the number of iterations. However, more generally, the analysis below holds whenever the number of samples within the LTMA loops does not exceed that of the rest of the algorithm (i.e. $\tilde{\mathcal{O}}(\epsilon^{-3})$ as will be shown below).
@@
Choose parameter settings according to Lemma 9, and let $\alpha=\frac{2\gamma^{2}m\eta_{\lambda}}{(1-\gamma)^{3}}$, $\eta_{\lambda}=1$, and $\eta=\frac{1-\gamma}{\alpha}$. Moreover, let $\epsilon_{k}^{\prime}>0$ be the upper bound on the LTMA error tolerance at macro-iteration $k$ such that $\epsilon_{k}^{\prime}=\Theta(\epsilon)$, and let $\mathcal{P}$ be an uncertainty set contained in an $\ell_{1}$ $(s,a)$-rectangular uncertainty set around the nominal distribution $\bar{p}=p_{0}$ such that
$\mathcal{P}_{s,a}\subseteq\left\{p\in\Delta(\mathcal{S}):\left\lVert p(\cdot|s,a)-\bar{p}(\cdot|s,a)\right\rVert_{1}\leq\psi_{s,a}\right\}\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}\,,$ (40)
and let $\Delta_{p}=\max_{s,a}\psi_{s,a}=\Theta\left(\frac{(1-\gamma)}{\alpha\log(A)}\right)$. Then within a total number of queries to the generative model equal to $T=\sum_{k=1}^{K}\left(M_{V,k}N_{V,k}+\sum_{t=0}^{t_{k}-1}M_{Q,k}N_{Q,k}+t_{k}^{\prime}\right)=\tilde{\mathcal{O}}(\epsilon^{-3})$, Robust Sample-based PMD-PD provides three guarantees.
@@
$\max_{j\in[m]}\frac{1}{K}\sum_{k=1}^{K}V_{\pi_{k},p_{k}}^{j}(\rho)\leq\mathcal{O}(\epsilon)\,.$ (42)
c) robust Lagrangian regret upper bound: under the conditions of Lemma 2, and with a number of LTMA iterations $t_{k}^{\prime}=\Theta\left(\frac{F_{\lambda_{k}}M}{\epsilon_{k}^{\prime}(1-\gamma}\right)$, we have

### v2

@@ 5.4 Analysis of Robust Sample-based PMD-PD
Note that
$\Phi(\pi_{k})$ $=\mathbf{V}_{\pi_{k},p_{k-1}^{*}}(\rho;\lambda_{k-1}^{*})$
$=\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi_{k},p}(\rho;\lambda_{k-1})+\left(\mathbf{V}_{\pi_{k},p_{k-1}^{*}}(\rho;\lambda_{k-1}^{*})-\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi_{k},p}(\rho;\lambda_{k-1})\right)$ (decomposing)
$\leq\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi_{k},p}(\rho;\lambda_{k-1})+\left(\mathbf{V}_{\pi_{k},p_{k-1}^{*}}(\rho;\lambda_{k-1}^{*})-\mathbf{V}_{\pi_{k},p_{k-1}^{*}}(\rho;\lambda_{k-1})\right)$ ($p_{k-1}^{*}$ is optimal for $\lambda_{k-1}^{*}$ but not for $\lambda_{k-1}$)
@@
$=\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi_{k},p}(\rho;\lambda_{k-1})+\left(V_{\pi_{k},p_{k-1}^{*}}(\rho)-V_{\pi_{k},p_{k-1}^{*}}(\rho)+\sum_{j=1}^{m}\left(\lambda_{k-1,j}^{*}-\lambda_{k-1,j}\right)V_{\pi_{k},p_{k-1}^{*}}^{j}(\rho)\right)$ (definition of Lagrangian, rearranging)
$\leq\mathbf{V}_{\pi_{k},p_{k}}(\rho;\lambda_{k-1})+\epsilon_{k}+\Delta_{\lambda_{k}}\,$,
where the last step follows from the error tolerance of LTMA and the definition of $\Delta_{\lambda_{k}}$.
Eq. 35 then follows from $\Phi(\pi^{*})=\sup_{p\in\mathcal{P}}\max_{\lambda\geq 0}\mathbf{V}_{\pi^{*},p}(\rho;\lambda)\geq\mathbf{V}_{\pi^{*},p_{k}}(\rho;\lambda_{k-1})$. ∎
@@
As sample-based algorithms work with approximate value functions, Lemma 9 describes the conditions under which one can obtain an $\epsilon$-approximation of the cost functions and the regularised augmented Lagrangian. The same settings as sample-based PMD in Lemma 15 of Liu et al., 2021b transfer to the Robust Sample-based PMD algorithm. We rephrase the lemma to include updated transition dynamics.
Lemma 9 (Value function approximation, Lemma 15 of Liu et al., 2021b rephrased).
@@
$\text{{a)}}\quad|\hat{V}_{\pi_{k}^{t},p_{k}}^{i}(\rho)-V_{\pi_{k}^{t},p_{k}}^{i}(\rho)|\leq\epsilon\quad\quad\forall i=1\in[k]$
$\text{{b)}}\quad|\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha}(s,a)-\tilde{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha}(s,a)|\leq\epsilon\quad\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}\,$,
@@
The full proof can be found in Appendix D.5. ∎
@@
Lemma 10 (Bound on regularised augmented Lagrangian under approximate entropy-regularised NPG, Lemma 18 in Liu et al., 2021b ).
@@
$\tilde{\mathbf{V}}_{\pi_{k+1},p_{k}}(s)+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k+1},p_{k}}}(\pi_{k+1},\pi_{k})\leq\tilde{\mathbf{V}}_{\pi,p_{k}}(s)+\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k})-B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k+1})\right)+\Theta(\epsilon)$ (36)
@@
The proof makes use of Lemma 19. In particular, it notes that $K=\Theta(1/\epsilon)$ and the value difference to $\pi_{k}^{*}$ is bounded by $1/K$ after $t_{k}$ steps. After applying the pushback property (Lemma 16) to $\pi$, $\pi_{k}$, and $\pi_{k}^{*}$, and derives an upper bound on the Bregman divergence w.r.t. $\pi_{k}^{*}$ based on $B_{d_{\rho}^{\pi,p_{k}}}(\pi,\pi_{k+1})+\left\lVert\log(\pi_{k}^{*})-\log(\pi_{k}^{t+1})\right\rVert_{\infty}$, where the second term is $\Theta(1/K)$. We refer to Lemma 18 of Liu et al., 2021b for the detailed proof. ∎
The theorem below proves the sample complexity of Robust Sample-based PMD-PD, using Lemma 9 to compute estimates of the Q-values for policy updates and using Lemma 6 to compute estimates of the G-values for transition kernel updates. This then allows to demonstrate the overall sample complexity for obtaining an average regret in the value, constraint-cost, and Lagrangian bounded by $\mathcal{O}(\epsilon)$.
@@
Choose parameter settings according to Lemma 9, let $\alpha=\frac{2\gamma^{2}m\eta_{\lambda}}{(1-\gamma)^{3}}$, $\eta_{\lambda}=1$, and $\eta=\frac{1-\gamma}{\alpha}$. Moreover, let $\pi^{*}\in\mathrm{arg}\inf_{\pi\in\Pi}\Phi(\pi)$, let $\epsilon_{k}^{\prime}>0$ be the upper bound on the LTMA error tolerance at macro-iteration $k$ such that $\epsilon_{k}^{\prime}=\Theta(\epsilon)$, and let $\mathcal{P}$ be an $\ell_{q}$ uncertainty set as in the precondition of Lemma 23. Then within a total number of queries to the generative model equal to $T=\tilde{\mathcal{O}}(\epsilon^{-3})$, Robust Sample-based PMD-PD provides the following guarantees:
@@
$\max_{j\in[m]}\frac{1}{K}\sum_{k=1}^{K}V_{\pi_{k},p_{k}}^{j}(\rho)=\mathcal{O}(\epsilon)\,.$ (38)
c) robust Lagrangian regret upper bound under LTMA oracle: under the conditions of Lemma 2, with a number of LTMA iterations $t_{k}^{\prime}=\Theta\left(\frac{F_{\lambda_{k}}M}{\epsilon_{k}^{\prime}(1-\gamma}\right)$, we have
d) robust Lagrangian regret upper bound under approximate LTMA: under the conditions of Lemma 7 as applied to the Lagrangian at each macro-iteration, with $C=F_{\lambda}$, we have